# From Empirical Observations to Universality: Dynamics of Deep Learning with Inputs Built on Gaussian mixture

### v1

@@ V.1.2 Dominance of Moments in the Expectation Value of Functions
Under this setting, for specific functions $f$ where higher-order differential terms are insignificant, the following lemma is derived:
@@
[…]
$\mathbb{E}_{x\sim\mathcal{P}_{\mathcal{D}2}}[f(x)]=\mathbb{E}_{x\sim\mathcal{D}2}[f(x)]$
The derivation hinges on the function $f$’s property, which erase the effect of high-order related terms, leading to convergence to the same expectation value. Detailed derivation can be found in the appendix C.
This suggests that for any random variable $x$ standardized to mean zero and variance one, the expected value of its function, $\mathbb{E}_{x\sim\mathcal{P}}[f(x)]$, can be closely approximated by $\mathbb{E}_{x\sim\mathcal{N}}[f(x)]$.
Thus, for a standardized Gaussian mixture distribution $\mathcal{P}$, since the nonlinear function $f\equiv\text{sgn}$ satisfies the condition, the expectations of covariance components in the Gaussian mixture align closely with their Gaussian counterparts:
$Q_{k,\ell}\equiv\mathbb{E}_{C\sim\mathcal{N}}\left[{\bar{\lambda}}_{k}{\bar{\lambda}}_{\ell}\right]\approx\mathbb{E}_{C\sim\mathcal{P}}\left[{\bar{\lambda}}_{k}{\bar{\lambda}}_{\ell}\right]$ (22)
$R_{k,m}\equiv\mathbb{E}_{C\sim\mathcal{N}}\left[\bar{\lambda}_{k}\nu_{m}\right]\approx\mathbb{E}_{C\sim\mathcal{P}}\left[\bar{\lambda}_{k}\nu_{m}\right]$ (23)
$T_{m,n}\equiv\mathbb{E}_{C\sim\mathcal{N}}\left[\nu_{m}\nu_{n}\right]\approx\mathbb{E}_{C\sim\mathcal{P}}\left[\nu_{m}\nu_{n}\right]$ (24)
The summary of above convergence can summarized under Modified Equivalence Property:
@@
With the same condition of II.3 except and the distribution of $C$ is standardized Gaussian mixture $\overline{\mathcal{P}}$, take the Gaussian Equivalence Property’s results $\{\lambda,\nu\}$ distribution as $\mathcal{G}$. Then the under standardized Gaussian mixture random variable $C$, current preactivation distribution $\{\lambda,\nu\}$ follows $\tilde{\mathcal{G}}$ such that
$\tilde{\mathcal{G}}\equiv\tilde{\mathcal{G}}_{\mathcal{G}_{2}}$ (25)
It is important to note that $\tilde{\mathcal{G}}$ is not follow Gaussian distribution but only shares identical cumulants with $\mathcal{G}$ up to order 2.
Surprisingly, if the neural network’s activation function $g$, $\tilde{g}$ also satisfy the condition for function expectation approximation V.3, results to their correlation with the random variables ${\lambda,\nu}$, such as $\mathbb{E}[g(\lambda_{k})\tilde{g}(\nu_{n})]$, used in the ODE formulation of $v$, also exhibits an approximated equivalence.
Fortunately, where the ReLU serves as the activation function, its piecewise second derivative is inherently zero. The higher order term in Taylor expansion is less dominant then the lower order term. Thus, a loose approximation is viable:
$\mathbb{E}_{\{\lambda,\nu\}\sim\mathcal{G}}\left[g(\lambda_{k})\tilde{g}(\nu_{n})\right]\approx\mathbb{E}_{\{\lambda,\nu\}\sim\tilde{\mathcal{G}}_{\mathcal{G}_{2}}}\left[g(\lambda_{k})\tilde{g}(\nu_{n})\right]$ (26)
Thus, even without higher order equivalence between $\mathcal{G}$ and $\tilde{\mathcal{G}}=\tilde{\mathcal{G}}_{\mathcal{G}_{2}}$, the core dynamics governing the neural network exhibit approximate equivalence.
In summary, our findings articulate the following points:
@@
    Specific functions $f$ that meet the criteria outlined, predominantly influence the function expectation by the first and second cumulants, V.3.
@@
    This adaptation and the conditions specified lead to an equivalence in the covariance of preactivations $\{\lambda,\nu\}$, achieve equivalence up to second cumulants from a distribution perspective, V.4.
@@
    The dynamics of the neural network, as dictated by function correlations involving the activation function and the random variables.
    V.4 and V.3 with activation function, yielding even dynamic equivalence under standardized Gaussian mixture.
@@
If erf were used as the activation function, since function is bounded $\in(-1,1)$, the conditions for lemma V.3 would have been satisfied more loosely, and a similar discussion could be applicable.
@@
Ultimately, despite the input distributions deviating from a Gaussian form, within the thermodynamic limit ($N\to\infty$, $D\to\infty$) and for specific functions that adhere to the conditions of lemma V.3—where the expectation value is dominated by the first and second moments—the dynamics of neural networks asymptotically converge to those anticipated under simple Gaussian inputs. This convergence facilitates the incorporation of dynamics observed under Gaussian mixtures into conventional theoretical frameworks originally devised for single Gaussian inputs.
This broadens the applicability of these theories, extending their relevance to encompass scenarios not previously considered in analyses confined to simple Gaussian inputs, thereby enhancing our theoretical understanding of neural networks across a more diverse array of input distributions.
@@ VI Conclusion
Previous research has extensively explored the dynamics under a hidden manifold with a simple Gaussian distribution [9, 10, 11], as well as under Gaussian mixtures in the absence of a hidden manifold [27, 28]. Our study represents a expansion of previous works, aiming to broaden the scope of understanding. We delved into investigating dynamics neural network when exposed to Gaussian mixtures in manifold—a more comprehensive representation of distributions. We scrutinized the discrepancies between dynamics under simple Gaussian based input and dynamics under Gaussian mixture based input with SGD update rule, yielding insightful findings.
The key takeaway from our investigation is the pivotal role of standardization. Applying standardization to Gaussian mixtures facilitates an alignment with the predictions posited by pre-existing conventional theories [10]. Furthermore, our rigorous mathematical substantiation enabled the extension of these conventional theories to accommodate a broader spectrum of distributions. This results underscores that, notwithstanding variations in the input structure from a Gaussian baseline, the dynamics of the student model essentially adhere to conventional theoretical models.
In essence, our study unveils a newfound universality in which the inferences drawn from function correlation and the Gaussian Equivalence Property retain their validity within the domain of Gaussian mixtures. Given the generalized Gaussian mixture model assumption and its substantial expressive capacity, it stands to reason that a similar universality might be observed across other distributions. Moreover, our mathematical justification, predicated on the significance of the first and second moments (V.3, V.4), paves the way for extending this framework to encompass a more diverse array of distributions, both traditional and beyond.
The inherent structural properties of data offer an intriguing and insightful foundation for exploration. By incorporating these characteristics into the analysis of deep learning dynamics, we aim to bridge the gap between the impressive practical successes of applied deep learning and the still-emerging theoretical research. Ultimately, we hope our research seeks to deepen the understanding of deep learning, contributing to further breakthroughs and insights in the field.
@@ VII Reproducibility
For a comprehensive understanding of our numerical SGD implementation and ODE update mechanisms, along with ensuring reproducibility, please visit our code repository at the following link: https://github.com/peardragon/_GaussianMixture.
@@ Acknowledgements.
Appendix: From Empirical Observations to Universality:
Dynamics of Deep Learning with Inputs Built on Gaussian mixture
A Derivation of Gaussian Equivalent Property
0.1 Correlation of Two Functions
@@
0.2 Gaussian Equivalence Property
@@
$\frac{dv_{k}}{dt}=\eta\left[\sum^{M}_{n}\tilde{v}_{n}\mathbb{E}[g(\lambda_{k})\tilde{g}(\nu_{n})]-\sum^{K}_{j}v_{j}\mathbb{E}[g(\lambda_{k}){g}(\lambda_{j})]\right]$ (S9)
Given the crucial role of function correlation in unpacking the dynamics prompted by weight updates, it is imperative to understand the distribution characterizing ${\lambda,\nu}$ to compute expectation values such as $\mathbb{E}[g(\lambda_{k})\tilde{g}(\nu_{n})]$. This analytical approach enables a deeper understanding of the underlying mechanics governing the behavior of neural networks, particularly in how weight adjustments influence overall learning and adaptation processes.
@@
In the thermodynamic limit $N\to\infty$, $D\to\infty$, matrices $W$, $\tilde{W}$, and $F$ possess explicit bounds:
@@
Even when considering matrices $F$ and $W$ together, they maintain explicit bounds. For all $p,q\geq 1$ and any indices $k_{1},\cdots,k_{p},r_{1},\cdots,r_{p}$:
@@
For convenience, let’s redefine $\bar{\lambda}_{k}$ as:
$\bar{\lambda}_{k}=\cfrac{1}{\sqrt{N}}\sum_{i=1}^{N}W_{k,i}(f(U_{i})-\mathbb{E}_{u\sim\mathcal{N}(0,1)}[f(u)])$ (S12)
$\bar{\lambda}_{k}$ also follows a jointly Gaussian distribution, and its expectation value satisfies $\mathbb{E}[\bar{\lambda}_{k}]=0$ as per function correlation.
@@
$Q_{k,\ell}$ $\equiv\mathbb{E}\left[{\bar{\lambda}}_{k}{\bar{\lambda}}_{\ell}\right]$ (S14)
@@
Consequently, the new distribution $\{\bar{\lambda},\nu\}$ follows a more straightforward distribution with the mean
$\mathbb{E}\left[\bar{\lambda}_{k}\right]=\mathbb{E}\left[\nu_{m}\right]=0$ (S21)
@@
$Q_{k,\ell}$ $\equiv\mathbb{E}\left[{\bar{\lambda}}_{k}{\bar{\lambda}}_{\ell}\right]=\left(c-a^{2}-b^{2}\right)\Omega_{k,\ell}+b^{2}\Sigma_{k,\ell}$ (S22)
$R_{k,m}$ $\equiv\mathbb{E}\left[\bar{\lambda}_{k}\nu_{m}\right]=b\frac{1}{D}\sum_{r=1}^{D}S_{k,r}\tilde{W}_{m,r}$ (S23)
$T_{m,n}$ $\equiv\mathbb{E}\left[\nu_{m}\nu_{n}\right]=\frac{1}{D}\sum_{r=1}^{D}\tilde{W}_{m,r}\tilde{W}_{n,r}.$ (S24)
@@
B Derivation of the ODE for Covariance and Weights

### v2

@@ V.1.2 Dominance of cumulants in the expectation values of functions
Under this setting, for specific functions $f$ where higher-order differential terms are insignificant, the following lemma is derived.
@@
Let $x$ be a random variable with mean $\mu$ and variance $\sigma^{2}$ under distribution $\mathcal{P}_{\mathcal{D}2}$. Suppose $f$ is a $C^{\infty}$ function almost everywhere, with the conditions that for $x\sim\mathcal{P}_{\mathcal{D}_{2}}$ and $x\sim\mathcal{D}$: (1) $\mathbb{E}[f(x)]_{x\in(\mu-\epsilon,\mu+\epsilon)}\to f(\mu)$, and (2) $(x-\mu)^{n}f^{(n)}(\mu)$ for $n>2$ is negligible in the particular limit of our interest. Then the function expectation possesses the following approximate property:
$\mathbb{E}_{x\sim\mathcal{P}_{\mathcal{D}_{2}}}[f(x)]=\mathbb{E}_{x\sim\mathcal{D}_{2}}[f(x)].$ (19)
The derivation hinges on the property of function $f$, which diminishes the effect of higher-order terms, leading to convergence to the same expectation value. A detailed derivation can be found in Appendix 1. This suggests that for any random variable $x$ standardized to mean 00 and variance $1$, the expected value of its function, $\mathbb{E}_{x\sim\mathcal{P}}[f(x)]$, can be closely approximated by $\mathbb{E}_{x\sim\mathcal{N}}[f(x)]$.
Thus, for a standardized Gaussian mixture distribution $\mathcal{P}$, since the nonlinear function $f\equiv\text{sgn}$ satisfies the above two conditions, the covariance matrices in the Gaussian mixture align closely with their Gaussian counterparts:
$Q_{k,\ell}\equiv\mathbb{E}_{C\sim\mathcal{N}}\left[{\overline{\lambda}}_{k}{\overline{\lambda}}_{\ell}\right]\approx\mathbb{E}_{C\sim\mathcal{P}}\left[{\overline{\lambda}}_{k}{\overline{\lambda}}_{\ell}\right],$ (20)
$R_{k,m}\equiv\mathbb{E}_{C\sim\mathcal{N}}\left[\overline{\lambda}_{k}\nu_{m}\right]\approx\mathbb{E}_{C\sim\mathcal{P}}\left[\overline{\lambda}_{k}\nu_{m}\right],$ (21)
$T_{m,n}\equiv\mathbb{E}_{C\sim\mathcal{N}}\left[\nu_{m}\nu_{n}\right]\approx\mathbb{E}_{C\sim\mathcal{P}}\left[\nu_{m}\nu_{n}\right].$ (22)
The above convergence can be summarized under the following modified equivalence property.
@@
With the same conditions in the GEP (Property II.3) and representing the preactivation distribution in GEP that follows jointly Gaussian as $\mathcal{G}$, i.e., $\{\lambda,\nu\}\sim\mathcal{G}$, the preactivation distribution under the standardized Gaussian mixture random variable $C\sim\overline{\mathcal{P}}$ follows $\widetilde{\mathcal{G}}$ such that
$\widetilde{\mathcal{G}}\equiv\widetilde{\mathcal{G}}_{\mathcal{G}_{2}}.$ (23)
It is important to note that $\widetilde{\mathcal{G}}$ does not follow a Gaussian distribution but only shares identical cumulants with $\mathcal{G}$ up to order 2.
Surprisingly, if the activation functions $g$ and $\widetilde{g}$ also satisfy the conditions for the approximation of the function expectation (Lemma V.3), their correlation with the random variables ${\lambda,\nu}$, such as $\mathbb{E}[g(\lambda_{k})\widetilde{g}(\nu_{n})]$, also exhibits an approximate equivalence.
For example, when ReLU serves as the activation function, its piecewise second derivative is 00. The higher-order term in the Taylor expansion is thus less dominant than the lower-order term, and therefore an approximation is viable:
$\mathbb{E}_{\{\lambda,\nu\}\sim\mathcal{G}}\left[g(\lambda_{k})\widetilde{g}(\nu_{n})\right]\approx\mathbb{E}_{\{\lambda,\nu\}\sim\widetilde{\mathcal{G}}_{\mathcal{G}_{2}}}\left[g(\lambda_{k})\widetilde{g}(\nu_{n})\right].$ (24)
Accordingly, even without higher-order equivalence between $\mathcal{G}$ and $\widetilde{\mathcal{G}}=\widetilde{\mathcal{G}}_{\mathcal{G}_{2}}$, the core dynamics governing the neural network in this work exhibit approximate equivalence.
In summary, our findings articulate the following points.
@@
    For specific functions $f$ that meet the criteria outlined, the function correlation is predominantly influenced by the first and second cumulants (Lemma V.3).
@@
    This adaptation and the specified conditions lead to an equivalence in the $\{\lambda,\nu\}$ distribution, achieving equivalence up to the second cumulants from a distribution perspective (Property V.4).
@@
    The dynamics of the neural network is dictated by the function correlation. Property V.4 and Lemma V.3, with higher-order Taylor terms diminished by the activation function, yield dynamic equivalence even under standardized Gaussian mixtures.
@@
If the erf function were used as the activation function, since the function is bounded as $\text{erf}(\cdot)\in(-1,1)$, the conditions for Lemma V.3 would be more loosely satisfied, and a similar discussion could be applicable.
V.2 Results from various distribution settings
As our mathematical proof shows, the only conditions for dynamics convergence are (1) $\mathcal{P}\equiv\mathcal{N}_{\mathcal{D}_{2}}$, and (2) diminishing higher-order derivatives of the nonlinear function $f$ and activation functions $g$ and $\widetilde{g}$. Therefore, for an arbitrary distribution with standardization, convergence results are expected to be obtained. We consider several distributions that are distinct from simple Gaussian and observe the SGD dynamics. For standardization, we take expectation values from samples to calculate sample mean and sample variance. The distribution parameter information is summarized in Table 1. For different nonlinear function $f$ settings, results under $f=tanh(\cdot)$ can be found in Appendix 2.
Table 1: Parameters for various distributions used in the experiments
Distribution Parameter(s)
uniform $a=0$, $b=10$
beta (1) $\alpha=0.5$, $\beta=0.5$
beta (2) $\alpha=5$, $\beta=1$
Poisson $\lambda=2$
Laplace $\mu=0$, $b=1$
Pareto $\alpha=5$
Lorentz $x_{0}=0$, $\gamma=1$
Gaussian mixture
$p_{i}=0.3,0.7$,
$\mu_{i}\sim\mathcal{U}[-2,2]$, $\sigma_{i}\sim\mathcal{U}[0.5,5]$
Figure 7: Examples of dynamics under various distribution settings. Dynamics of (a) generalization error $\epsilon_{g}$, (b) covariance matrix $Q$, (c) covariance matrix $R$, and (d) weight of the second layer $v$. The SGD results are averaged over 10 runs.
Figure 8: Example of $v_{k}$ dynamics under various distribution settings, including the Lorentz distribution.
In Fig. 7, we observe the expected convergence phenomena. Additionally, if we use a distribution that has no low-order cumulant, standardization cannot satisfy the $\mathcal{P}\equiv\mathcal{N}_{\mathcal{D}_{2}}$ assumption theoretically. To strengthen our discussion through this example, we consider the Lorentz distribution, which has no mean or variance. For compact visualization, we use the dynamics of $v_{k}$.
@@ V.2 Results from various distribution settings
In Fig. 8, we see that even standardization cannot make the dynamics converge. This direct example reaffirms that the cumulant property is the key factor for universality. Ultimately, despite the distinct input distributions deviating from simple Gaussian, standardization and the specific functions that adhere to the conditions of Lemma V.3—where the expectation values are dominated by the first and second cumulants—lead the dynamics of the neural network to asymptotically converge to those anticipated under simple Gaussian inputs.
V.3 Application on Pseudo-Real Data
Figure 9: Samples of pseudo-real dataset $\hat{X}=f(\overline{C}F/\sqrt{D})$ (first and third rows) and true MNIST dataset $X_{\text{MNIST}}$ (second and fourth rows).
Figure 10: Examples of dynamics under the pseudo-real MNIST dataset. SGD results are plotted as dashed lines. Both SGD and ODE results are averaged over 10 runs.
The stringent assumptions of a random teacher and random dataset ($C$) limit the scope of this analytical study. By extending the analysis under low-order cumulant dominance, we can explore further possibilities.
The main premise here is to account for weak correlations rather than treating variables as strictly independent. In practical settings, it is unrealistic to generate fully independent datasets, and thus some degree of generalization can be achieved by acknowledging weak correlations. To construct a pseudo-real dataset within this framework, three components are required—$C$, $F$, and the teacher model. The first two components ($C$ and $F$) serve as inputs for the teacher and student models. Under standardization, $C\sim\overline{\mathcal{P}}$, or $\overline{C}$, the student input $X=f(\overline{C}F/\sqrt{D})=f(U)$ should resemble a real dataset, with $\overline{C}\in\mathbb{R}^{D}$ having a lower dimension than the actual dataset $X\in\mathbb{R}^{N}$. Additionally, the teacher model must establish a relationship with the true labels, $y$. In this study, we mimic the MNIST dataset.
V.3.1 Modeling the pseudo-real dataset
Before proceeding, it is important to recheck the vital assumptions that constrain our analytical descriptions. Since we extend the framework for $C$ to cover an arbitrary distribution, we need to consider only the correlation of $U=\overline{C}F/\sqrt{D}$. In our analytical assumption, $C$ needs to have properties resulting in $\mathbb{E}[U_{i}U_{i}]\approx 1$ and weak correlations $\mathbb{E}[U_{i}U_{j}]\approx\epsilon$ where $i\neq j$.
Since our goal is to replicate a real dataset, $C$ and $F$ must be related such that $X=f(\overline{C}F/\sqrt{D})$ closely approximates the MNIST dataset. This problem can be formulated as an inverse optimization task to find a suitable $C$ such that $\hat{X}=f(U)$ not only minimizes the discrepancy between $X_{\text{MNIST}}$ and $\hat{X}$ but also preserves the weak correlation $\mathbb{E}[U_{i}U_{j}]\approx\epsilon$. This combined consideration can be represented as an optimization problem of $C$ with a modified loss function $\mathcal{L}_{C}$:
$\mathcal{L}_{C}=|\hat{X}-X_{\text{MNIST}}|+|\mathbb{E}[U_{i}U_{j}]-\mathbb{E}_{u\sim\mathcal{N}(0,I)}[u_{i}u_{j}]|.$ (25)
We take $10^{4}$ samples from the MNIST dataset, $X_{\text{MNIST}}\in\mathbb{R}^{10^{4}\times 28^{2}}$, and generate $F$ by element-wise Gaussian entries to satisfy the bounded assumption (Assumption II.2). By incorporating these constraints into a modified loss function and applying a quasi-Newton method (the limited-memory BFGS optimizer), we determine $C$ with dimension $D=500$ that both mimics the MNIST dataset and maintains numerically weak correlations. We plot examples of $\hat{X}$ generated under optimized $C$ and $X_{\text{MNIST}}$ in Fig. 9. Additionally, since we need a specific distribution $\mathcal{P}$ to sample $C$, we use kernel density estimation (KDE) for the probability density estimation.
V.3.2 Modeling the pseudo–teacher model
Next, we need to set the pseudo–teacher model, referring to a model that effectively knows the true labels. This requires training the teacher model to predict the correct labels, for which we use a binary labeling of the MNIST dataset, classifying digits as odd or even. The teacher model employs a $hardtanh$ activation function, $hardtanh(x)=\max(-1,\min(1,x))$, to output values indicating whether the digit is more odd-like ($+1$) or even-like ($-1$).
A key assumption at this stage is that the teacher model’s weights exhibit weak correlations both within the model and with the input data. In detail, it is essential to preserve the correlations $\mathbb{E}[\nu_{i}^{2}]$ and $\mathbb{E}[\nu_{i}\nu_{j}]$ where $\nu=C\widetilde{W}^{\top}/\sqrt{D}$ as much as possible during training. The optimization task here is to ensure that the teacher model’s output $\hat{y}$ matches the binary MNIST labels $y_{\text{MNIST}}$ while preserving the correlation. In summary, this stage involves the optimization problem of the teacher model with a modified loss function:
$\mathcal{L}_{\text{teacher}}=|\hat{y}-y_{\text{MNIST}}|+|\mathbb{E}[\nu_{i}\nu_{j}]-\mathbb{E}_{\text{Initial}}[\nu_{i}\nu_{j}]|.$ (26)
As expected, a trade-off arises between correlation preservation and target accuracy. Nevertheless, because the teacher model has vast parameters, it is possible to numerically preserve the correlations while achieving meaningful results. With the modified loss function and limited-memory BFGS optimizer, we obtain a pseudo–teacher model with a $65.58\%$ accuracy in predicting the true labels, while mostly preserving the correlations.
This accuracy is statistically significant based on a t-test comparison with randomly initialized teacher models, suggesting that the pseudo–teacher model effectively captures and utilizes the input dataset’s characteristics for labeling.
V.3.3 Pseudo-real dataset results
With the optimized $C$ and pseudo–teacher model, we obtain the SGD dynamics as in previous sections. Before comparing these dynamics with analytical predictions, we need a few corrections to the equations as there are differences between the numerically weak correlation and analytically independent settings. The expectation values related to $U=CF/\sqrt{D}$ can no longer be considered under $U$ from a normal Gaussian distribution. We corrected the expectation values as follows: $a=\mathbb{E}[f(U)]$, $b=\mathbb{E}[Uf(U)]$, and $c=\mathbb{E}[f(U)^{2}]$.
With the same initial order parameters $Q(t=0)$ and $R(t=0)$, the analytical predictions are calculated by ODEs as in previous sections. As illustrated in Fig. 10, even with the pseudo-real dataset, the system dynamics are well preserved and exhibit asymptotic behavior consistent with the analytical predictions. This extension to allow arbitrary distributions as inputs offers a foundation for exploring deep learning behaviors in real-world domains. Similar results with different teacher models but comparable accuracy can be found in Appendix 3.
@@ VI Conclusion
Previous studies have examined the dynamics of neural networks under simple Gaussian distributions and Gaussian mixtures without hidden manifolds [18, 19, 20, 38, 39, 40]. Our work extends this by investigating dynamics with Gaussian mixtures in a manifold setting. We compared dynamics under simple Gaussian inputs to those under Gaussian mixture inputs using SGD, identifying key differences.
The key takeaway from our study is the pivotal role of standardization when using functions whose high-order differential terms diminish. Applying standardization to any distribution with existing cumulants facilitates an alignment with the dynamics observed under a simple Gaussian distribution. Standardization also allows for convergence of dynamics even with general distributions and when using conventional activation functions such as ReLU and sigmoidal. In essence, these results may point to universality.
Numerical experiments with a pseudo-real dataset showed approximate alignment between our theoretical framework and observed behaviors, under approximate satisfaction of the weak correlation assumption. Future work could extend this framework by applying the findings to real-world datasets and relaxing some of the strict assumptions. Additionally, investigating how a dataset’s structural property (e.g., higher-order cumulants) influences the dynamics can broaden our understanding.
Exploring how deep learning parameters behave can be key to understanding the generalization ability of deep learning. We expect that our research will contribute to bridging the gap between the practical successes of applied deep learning and its developing theoretical foundations.
@@ VII Reproducibility
For a comprehensive understanding of our numerical SGD implementation and ODE update mechanisms and to ensure reproducibility, please visit our code repository at https://github.com/peardragon/GaussianUniversality.
@@ Acknowledgements.
Appendix: Gaussian Universality in Neural Networks Dynamics with Generalized Structured Input Distributions
A Background: Derivation of Gaussian Equivalent Property and ODE
1 Correlation of Two Functions
@@
2 Gaussian Equivalence Property
@@
$\frac{dv_{k}}{dt}=\eta\left[\sum^{M}_{n}\widetilde{v}_{n}\mathbb{E}[g(\lambda_{k})\widetilde{g}(\nu_{n})]-\sum^{K}_{j}v_{j}\mathbb{E}[g(\lambda_{k}){g}(\lambda_{j})]\right]$ (S9)
Given the crucial role of function correlation in unpacking the dynamics prompted by weight updates, it is imperative to understand the distribution characterizing ${\lambda,\nu}$ to compute expectation values such as $\mathbb{E}[g(\lambda_{k})\widetilde{g}(\nu_{n})]$. This analytical approach enables a deeper understanding of the underlying mechanics governing the behavior of neural networks, particularly in how weight adjustments influence overall learning and adaptation processes.
@@
In the thermodynamic limit $N\to\infty$, $D\to\infty$, matrices $W$, $\widetilde{W}$, and $F$ possess explicit bounds:
@@
Even when considering matrices $F$ and $W$ together, they maintain explicit bounds. For all $p,q\geq 1$ and any indices $k_{1},\cdots,k_{p},r_{1},\cdots,r_{q}$:
@@
For convenience, let’s redefine $\overline{\lambda}_{k}$ as:
$\overline{\lambda}_{k}=\cfrac{1}{\sqrt{N}}\sum_{i=1}^{N}W_{k,i}(f(U_{i})-\mathbb{E}_{u\sim\mathcal{N}(0,1)}[f(u)])$ (S12)
$\overline{\lambda}_{k}$ also follows a jointly Gaussian distribution, and its expectation value satisfies $\mathbb{E}[\overline{\lambda}_{k}]=0$ as per function correlation.
@@
$Q_{k,\ell}$ $\equiv\mathbb{E}\left[{\overline{\lambda}}_{k}{\overline{\lambda}}_{\ell}\right]$ (S14)
@@
Consequently, the new distribution $\{\overline{\lambda},\nu\}$ follows a more straightforward distribution with the mean
$\mathbb{E}\left[\overline{\lambda}_{k}\right]=\mathbb{E}\left[\nu_{m}\right]=0$ (S21)
@@
$Q_{k,\ell}$ $\equiv\mathbb{E}\left[{\overline{\lambda}}_{k}{\overline{\lambda}}_{\ell}\right]=\left(c-a^{2}-b^{2}\right)\Omega_{k,\ell}+b^{2}\Sigma_{k,\ell}$ (S22)
$R_{k,m}$ $\equiv\mathbb{E}\left[\overline{\lambda}_{k}\nu_{m}\right]=b\frac{1}{D}\sum_{r=1}^{D}S_{k,r}\widetilde{W}_{m,r}$ (S23)
$T_{m,n}$ $\equiv\mathbb{E}\left[\nu_{m}\nu_{n}\right]=\frac{1}{D}\sum_{r=1}^{D}\widetilde{W}_{m,r}\widetilde{W}_{n,r}.$ (S24)
@@
3 Derivation of the ODE for Covariance and Weights